# Deploying Open-Source Large Language Models: A Performance Analysis

### v1

@@ Abstract
Since the release of ChatGPT [5] in November 2023, large language models (LLMs) [1] have seen considerable success, including in the open-source community, with many open-weight models available. However, the requirements to deploy such a service are often unknown and difficult to evaluate in advance. To facilitate this process, we conducted numerous tests at the Centre Inria de l’Université de Bordeaux. In this article, we propose a comparison of the performance of several models of different sizes (mainly Mistral [2][3] and LLaMa [4]) depending on the available GPUs, using vLLM [8], a Python library designed to optimize the inference of these models. Our results provide valuable information for private and public groups wishing to deploy LLMs, allowing them to evaluate the performance of different models based on their available hardware. This study thus contributes to facilitating the adoption and use of these large language models in various application domains.
@@ 1 Introduction
Following the release of ChatGPT by OpenAI in November 2023 [5], large language models (LLMs) [1] have sparked great interest in the private sector, with many companies seeking to offer services based on these models. However, training and inference of such models remain inaccessible to the general public, requiring considerable computational power and high-quality data. For example, Meta AI acquired 350,000 NVIDIA H100 GPUs in January 2024, at an estimated cost of \$9 billion, specifically for training LLMs. Their training of LLaMa-3 was performed on $14*10^{12}$ tokens.

### v2

@@ Abstract
Since the release of ChatGPT [5] in November 2022, large language models (LLMs) [1] have seen considerable success, including in the open-source community, with many open-weight models available. However, the requirements to deploy such a service are often unknown and difficult to evaluate in advance. To facilitate this process, we conducted numerous tests at the Centre Inria de l’Université de Bordeaux. In this article, we propose a comparison of the performance of several models of different sizes (mainly Mistral [2][3] and LLaMa [4]) depending on the available GPUs, using vLLM [8], a Python library designed to optimize the inference of these models. Our results provide valuable information for private and public groups wishing to deploy LLMs, allowing them to evaluate the performance of different models based on their available hardware. This study thus contributes to facilitating the adoption and use of these large language models in various application domains.
@@ 1 Introduction
Following the release of ChatGPT by OpenAI in November 2022 [5], large language models (LLMs) [1] have sparked great interest in the private sector, with many companies seeking to offer services based on these models. However, training and inference of such models remain inaccessible to the general public, requiring considerable computational power and high-quality data. For example, Meta AI acquired 350,000 NVIDIA H100 GPUs in January 2024, at an estimated cost of \$9 billion, specifically for training LLMs. Their training of LLaMa-3 was performed on $14*10^{12}$ tokens.